OpenStudy (anonymous):

r/-3s, for r=-1/8 and s =3/4

OpenStudy (anonymous):

1/18

OpenStudy (anonymous):

Impossible8 - How did you solve the problem

OpenStudy (anonymous):

-3(3/4) = -(9/4) then -(1/8) / -(9/4) = (1/8) * (4/9) = (1/18)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

no problem =)